3) Production in the country of ECNLand can be characterized by the following function:
Y = 2K
0.2L
0.8 and that ECNLand initially has 200 units of labor and 800 units of capital.
a) Calculate the numerical value` marginal products of labor and marginal product of capital
1
Expert's answer
2018-10-12T15:52:09-0400

The numerical values of marginal products of labor and marginal product of capital are:

MPL = Y'(L) = 1.6K^0.2L^(-0.2) = 1.6*800^0.2*200^(-0.2) = 2.11,

MPK = Y'(K) = 0.4K^(-0.8)L^0.8 = 0.4*800^(-0.8)*200^0.8 = 0.13.